# DRT-o1: Optimized Deep Reasoning Translation via Long Chain-of-Thought

### v2

@@ Abstract
Recently, O1-like models have emerged as representative examples, illustrating the effectiveness of long chain-of-thought (CoT) in reasoning tasks such as math and coding tasks. In this paper, we introduce DRT-o1, an attempt to bring the success of long CoT to neural machine translation (MT). Specifically, in view of the literature books that might involve similes and metaphors, translating these texts to a target language is very difficult in practice due to cultural differences. In such cases, literal translation often fails to convey the intended meaning effectively. Even for professional human translators, considerable thought must be given to preserving semantics throughout the translation process. To simulate LLMs’ long thought ability in MT, we first mine sentences containing similes or metaphors from existing literature books, and then develop a multi-agent framework to translate these sentences via long thought. In the multi-agent framework, a translator is used to iteratively translate the source sentence under the suggestions provided by an advisor. To ensure the effectiveness of the long thoughts, an evaluator is also employed to quantify the translation in each round. In this way, we collect tens of thousands of long-thought MT data, which is used to train our DRT-o1. Using Qwen2.5 and LLama-3.1 as the backbones, DRT-o1 models can learn the thought process during machine translation, and outperform vanilla LLMs as well as existing O1-like LLMs, showing their effectiveness
DRT-o1: Optimized Deep Reasoning Translation via Long Chain-of-Thought
Jiaan Wang,  Fandong Mengthanks:   Corresponding author.,  Yunlong Liang,  Jie Zhou Pattern Recognition Center, WeChat AI, Tencent Inc {torchwang,fandongmeng,yunlonliang,withtomzhou}@tencent.com [Uncaptioned image]DRT-o1[Uncaptioned image]DRT-o1-7B[Uncaptioned image]DRT-o1-8B[Uncaptioned image]DRT-o1-14B
@@ 1 Introduction
Recently, the emergence of the O1-like models shows great performance in reasoning tasks, e.g., math and coding tasks OpenAI (2024b); Qin et al. (2024); Huang et al. (2024); Zhang et al. (2024); Zhao et al. (2024). With the help of long thought, LLMs tend to explore, reflect and self-improve the reasoning processes to achieve more accurate answers.
In this paper, our goal is not to achieve competitive performance with OpenAI’s O1 in neural machine translation (MT). Instead, we explore technical routes to bring the success of long thought to MT. To this end, we introduce DRT-o1, a byproduct of our exploration, and we hope it could facilitate the corresponding research in this direction. There are two key points in achieving this goal:
  1. i)
    A suitable translation scenario to employ long thought in MT: Not all scenarios require long thought during translation. For example, in simple expressions, literal translation can meet most needs, and translation via long thought may be unnecessary.
  2. ii)
    A method to synthesize MT data with long thought: Long thought SFT (supervised fine-tuning) data plays a vital role in simulating LLMs’ long thought ability Huang et al. (2024). Previous work typically pays much attention to how to synthesize long-thought SFT data in math and coding tasks Qin et al. (2024); Huang et al. (2024); Zhao et al. (2024).
@@
Based on the collected long-thought MT samples, we train (SFT) our DRT-o1-7B, DRT-o1-8B and DRT-o1-14B using the backbones of Qwen2.5-7B-Instruct, Llama-3.1-8B-Instruct Dubey et al. (2024) and Qwen2.5-14B-Instruct Qwen Team (2024), respectively. Experimental results on literature translation demonstrate the effectiveness of the DRT-o1. For example, DRT-o1-14B outperforms vanilla Qwen2.5-14B-Instruct and QwQ-32B-preview in terms of BLEU, CometKiwi, CometScore and GPT-4 evaluators.
@@
  • We propose DRT-o1 aiming at building LLMs with long-thought machine translation ability. To achieve this, we mine literature sentences with similes or metaphors, and collect MT samples with long-thought processes.
  • To synthesize the long-thought MT samples, we propose a multi-agent framework that involves a translator, an advisor and an evaluator. These three agents collaborate in an iterative manner to produce long thoughts during MT. Lastly, GPT-4o is used to further improve the quality of the synthetic long-thought MT samples.
  • Experimental results on literature translation verify the effectiveness of our DRT-o1. With the help of long thought, LLMs can learn to think during the machine translation.
2 DRT-o1 Data
We focus on English-to-Chinese translation, and we introduce how to collect DRT-o1 training data via three steps in this section: (1) collecting English sentences that tend to require long thoughts during translation (§ 2.1); (2) synthesizing the long-thought translation process for the collected sentences by a designed multi-agent framework (§ 2.2); (3) improving the readability and fluency of the long-thought content to form the final long-thought MT samples (§ 2.3). In the end, we provide data statistics of the collected data to give a deeper understanding (§ 2.4).
@@ 2.1 Literature Book Mining
First, we extract all sentences from these books, and filter out too short or too long sentences, i.e., less than 10 words or more than 100 words, resulting in 577.6K literature sentences. Second, for each sentence, we use Qwen2.5-72B-Instruct Qwen Team (2024) to judge whether the sentence involves similes or metaphors, and discard the sentences that do not contain any ones. Third, for the remaining sentences, we let Qwen2.5-72B-Instruct literally translate them to Chinese, and then judge whether the translation satisfies native Chinese people. If the answer is negative, the corresponding sentence will be reserved, and regarded as “suitable to translate via long thought”. In this manner, we finally collect 63K (out of 577.6K) literature sentences involving similes or metaphors whose literal translations also have flaws, called pre-collected sentences.
@@ 2.2 Multi-Agent Framework
For each pre-collected sentence (denoted as $s$), we design a multi-agent framework to translate it from English to Chinese via long thought. As shown in Figure 1, our multi-agent framework includes three agents: a translator, an advisor, and an evaluator. The synthetic process is illustrated as follows:
@@
(3) Translation Refine Loop. In the refine loop, three agents work together to refine the translation iteratively. In each iteration step $k$ (start from $k=1$), the advisor first evaluates the translation in the previous step, i.e., $t^{k-1}$, and provides detailed feedback $f^{k-1}$ for polishing it. Then, the evaluator gives an overall score of $t^{k-1}$ conditioned on both pre-defined scoring criteria and $f^{k-1}$, the score is denoted as $s^{k-1}$. In the last of the iteration step, the translator takes its previous translation $t^{k-1}$, the corresponding feedback $f^{k-1}$ and overall score $s^{k-1}$ into account to provide a new translation $t^{k}$. The translation refine loop will stop when the overall score reaches a pre-defined threshold or the number of iteration steps meets the maximum.
@@ 2.3 Long Thought Reformulation
where $1\leq r_{1}<r_{2}<...<r_{n}\leq m$, and $n$ is the number of remaining steps. If $n<3$, we will discard the whole sample, i.e., $\mathcal{P}(s)$ / $\mathcal{P}^{{}^{\prime}}(s)$.
For the remaining samples, we follow Qin et al. (2024), and leverage GPT-4o OpenAI (2024a) to modify and polish $\mathcal{P}^{{}^{\prime}}(s)$ into a self-reflection description. Finally, we obtain 22,264 machine translation samples with long thought. Figure 2 gives an example sample to illustrate the synthetic results.
@@
DRT-o1 data ${}_{\text{(training)}}$ 19,264 37.25 527.64 44.67
DRT-o1 data ${}_{\text{(validation)}}$ 1,000 37.43 531.36 44.98
DRT-o1 data ${}_{\text{(testing)}}$ 2,000 37.19 525.44 44.70
Table 1: The number of samples and average token-level length of query, thought and output. “Query” and “Output” in DRT-o1 data mean the source sentences and the translated outputs, respectively.
@@
Figure 3: The distribution of the number of refinement steps in DRT-o1 data.
2.4 Data Statistics
We split the collected 22,264 samples into training, validation and testing sets with 19,264, 1,000 and 2,000 samples, respectively. Table 1 shows the data statistics of DRT-o1 data and previous O1-like data. For Marco-O1 CoT data Zhao et al. (2024), since it is not fully released, we use its demo data to calculate the data statistics. [^2] https://github.com/AIDC-AI/Marco-o1 As we can see, the average number of tokens in our synthesized thought reaches 500+ tokens, showing the long thought process in our data.
Figure 3 shows the number of refinement steps/rounds in the DRT-o1 data, which ranges from three to eight. We can find that most samples (73.22%) involve three refinement steps, while only one sample involves eight steps. Furthermore, to provide a deeper understanding of the refinement process, we calculate the average edit distance before and after each refinement step. Specifically, the first three refinement steps cause 21.44, 13.16 and 10.90 character-level edit distance. This observation is consistent with intuition. As the refinement progresses, the magnitude of the modification gradually decreases.
@@ 2.4 Data Statistics
GRF CometKiwi GRB BLEU CometScore
@@
Llama-3.1-8B-Instruct 79.25 70.14 73.30 18.55 74.58
Qwen2.5-7B-Instruct 81.53 70.36 77.92 27.02 76.78
Qwen2.5-14B-Instruct 84.74 72.01 80.85 30.23 78.84
Marco-o1-7B 82.41 71.62 77.50 29.48 77.41
QwQ-32B-preview 86.31 71.48 83.08 27.46 78.68
DRT-o1
DRT-o1-8B (Backbone: ${}_{\text{(Backbone: Llama-3.1-8B-Instruct)}}$ 84.49 70.85 80.80 32.67 78.81
DRT-o1-7B (Backbone: ${}_{\text{(Backbone: Qwen2.5-7B-Instruct)}}$ 85.57 71.78 82.38 35.54 80.19
DRT-o1-14B (Backbone: ${}_{\text{(Backbone: Qwen2.5-14B-Instruct)}}$ 87.19 72.11 83.20 36.46 80.64
Table 2: Experimental results on literature translation.
@@ 3.1 Experimental Setups
Metrics. Following previous work, we adopt BLEU Papineni et al. (2002), reference-free CometKiwi and reference-based CometScore Rei et al. (2022) to evaluate the model translations. BLEU evaluates n-grams overlap between the generated translations and corresponding references, while CometScore evaluates the semantic similarity of translations against references. CometKiwi uses a language model to judge whether a translation conveys the semantics of the source sentence. To calculate CometKiwi and CometScore, we leverage the official codes [^3] https://github.com/Unbabel/COMET and the official models [^4] https://huggingface.co/Unbabel/wmt22-cometkiwi-da and https://huggingface.co/Unbabel/wmt22-comet-da. To calculate the BLEU score, we use the sacrebleu toolkit [^5] https://github.com/mjpost/sacrebleu to calculate the corpus-level BLEU.
@@
In addition, recent studies Kocmi and Federmann (2023); Wang et al. (2023) also show the strong ability of LLMs in NLP evaluation. Thus, we use evaluators implemented using GPT-4o in reference-based and reference-free styles, which we refer to as GRB and GRF, respectively. Since GRB and GRF need the API costs, we randomly select 400 samples from the testing set, and conduct the GRB and GRF evaluation. The evaluation prompts are given in Appendix A.
@@
Backbones. We adopt the following three LLMs as the backbones of our DRT-o1: (1) Llama-3.1-8B-Instruct Dubey et al. (2024) [^6] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct; (2) Qwen2.5-7B-Instruct [^7] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct and (3) Qwen2.5-14B-Instruct Yang et al. (2024) [^8] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct.
Implementation Details. Llama-Factory Zheng et al. (2024) is used to instruct-tune LLMs. Following Wang et al. (2024a), all LLMs are tuned on 8$\times$×NVIDIA A100 GPUs (40G) with 1e-5 learning rate and 8 (8$\times$1) batch size. We use the DeepSpeed optimization Rasley et al. (2020), and set ZeRO-3 optimization. Following Qin et al. (2024), we set the number of training epochs to 3, and the training process costs 70 GPU hours and 124 GPU hours for 7B and 14B models, respectively.
Inference Details. When evaluating model performance on the test set, we use vLLM toolkit Kwon et al. (2023) [^9] https://github.com/vllm-project/vllm to accelerate the model generation. We use the sampling decoding strategy with 0.1 temperature, and set the repetition penalty to 1.05.
Source Sentence Qwen2.5-14B-Instruct QwQ-32B-Preview DRT-o1-14B
Women were weeping and children crying, and all were going as fast as seemingly lay in their power, looking behind now and then as if pursued by some deadly enemy. {CJK}UTF8gbsn 妇女在哭泣,孩子们在哀号,所有人都似乎在尽全力快速前行,不时回头张望,仿佛背后有某个致命的敌人在追赶。 {CJK}UTF8gbsn 妇女们哭泣着,孩子们哭喊着,每个人都在尽力加快脚步,不时回头张望,仿佛背后有致命的敌人在追击。 {CJK}UTF8gbsn 妇女悲泣孩童哀号众人疾奔不时回首,仿佛有致命之敌在后紧追。
Wait through this season of the rain Till suns of autumn dry the plain, Then shall thy giant foe, and all His host and realm, before thee fall. {CJK}UTF8gbsn 等待这雨季过去,直到秋日的阳光晒干原野,那时你的巨敌及其部众和王国都将倒在你面前。 {CJK}UTF8gbsn 忍过雨季漫长夜待秋阳升空晒干大地湿润色,那时,巨敌垮台,他的军队和王国,全在你脚下 {CJK}UTF8gbsn 历尽这雨季,待秋日阳光驱散平原上的最后一丝湿润,那时,你那巨敌,连同他的军队与王国,都将匍匐在你脚下

### v3

# DRT: Deep Reasoning Translation via Long Chain-of-Thought

@@ Abstract
Recently, O1-like models have emerged as representative examples, illustrating the effectiveness of long chain-of-thought (CoT) in reasoning tasks such as math and coding tasks. In this paper, we introduce DRT, an attempt to bring the success of long CoT to neural machine translation (MT). Specifically, in view of the literature books that might involve similes and metaphors, translating these texts to a target language is very difficult in practice due to cultural differences. In such cases, literal translation often fails to convey the intended meaning effectively. Even for professional human translators, considerable thought must be given to preserving semantics throughout the translation process. To simulate LLMs’ long thought ability in MT, we first mine sentences containing similes or metaphors from existing literature books, and then develop a multi-agent framework to translate these sentences via long thought. In the multi-agent framework, a translator is used to iteratively translate the source sentence under the suggestions provided by an advisor. To ensure the effectiveness of the long thoughts, an evaluator is also employed to quantify the translation quality in each round. In this way, we collect tens of thousands of long-thought MT data, which is used to train our DRT. Using Qwen2.5 and LLama-3.1 as the backbones, DRT models can learn the thought process during machine translation, and outperform vanilla LLMs as well as LLMs which are simply fine-tuning on the paired sentences without long thought, showing its effectiveness.
DRT: Deep Reasoning Translation via Long Chain-of-Thought
Jiaan Wang,  Fandong Mengthanks:   Corresponding author.,  Yunlong Liang,  Jie Zhou Pattern Recognition Center, WeChat AI, Tencent Inc {torchwang,fandongmeng,yunlonliang,withtomzhou}@tencent.com [Uncaptioned image]DRT[Uncaptioned image]DRT-7B[Uncaptioned image]DRT-8B[Uncaptioned image]DRT-14B
@@ 1 Introduction
Recently, the emergence of the O1-like LLMs shows great performance in reasoning tasks, e.g., math and coding tasks OpenAI (2024b); Qin et al. (2024); Huang et al. (2024); Zhang et al. (2024); Zhao et al. (2024). With the help of long thought, LLMs tend to explore, reflect and self-improve the reasoning processes to achieve more accurate answers.
In this paper, we explore technical routes to bring the success of long thought to MT. To this end, we introduce DRT, a product of our exploration, and we hope it could facilitate the research community. There are two key points in achieving this goal:
i) A suitable translation scenario to employ long thought in MT: Not all scenarios require long chain-of-thought (CoT)111“long CoT” is equal to “long thought”, and we alternatively use these two terms in this paper. during translation. For example, in simple expressions, literal translation can meet most needs, and translation via long CoT may be unnecessary. Inappropriate scenarios might cause the overthinking issue Chen et al. (2024).
ii) A method to synthesize MT data with long thought: Long thought SFT (supervised fine-tuning) data plays a vital role in simulating LLMs’ long thought ability Huang et al. (2024). Previous work pays much attention to how to synthesize long-thought data in math and coding tasks Qin et al. (2024); Huang et al. (2024); Zhao et al. (2024).
@@
Based on the collected long-thought MT samples, we train our DRT-7B, DRT-8B and DRT-14B using the backbones of Qwen2.5-7B-Instruct, Llama-3.1-8B-Instruct Dubey et al. (2024) and Qwen2.5-14B-Instruct Yang et al. (2024a), respectively. Experimental results on literature translation verify their effectiveness. In particular, DRT-14B outperforms QwQ-32B-preview and DeepSeek-R1-Distill-Qwen-32B in terms of BLEU, CometKiwi, CometScore and GPT-4 evaluations. Moreover, human evaluation and case study show the strong translation performance of DRT models.
@@
  • We propose DRT aiming at building LLMs with long-thought machine translation ability. To achieve this, we mine literature sentences with similes or metaphors, and collect MT samples with long-thought processes.
  • To synthesize the long-thought MT samples, we propose a multi-agent framework that involves a translator, an advisor and an evaluator. These three agents collaborate in an iterative manner to produce long thoughts during MT. Lastly, GPT-4o is used to further improve the quality of the synthesized long-thought MT samples.
  • Experimental results on literature translation verify the effectiveness of our DRT. With the help of long thought, LLMs can learn to think during the machine translation.
2 DRT Data
We focus on English-to-Chinese translation [^2] Although we focus on English-to-Chinese translation in this work, the methods we introduced can be trivially applied to other languages or translation directions., and we introduce how to collect the long-thought MT samples via three steps in this section: (1) collecting English sentences that tend to require long thoughts during translation (§ 2.1); (2) synthesizing the long-thought translation process for the collected sentences by a designed multi-agent framework (§ 2.2); (3) improving the readability and fluency of the long-thought content to form the final long-thought MT samples (§ 2.3). Next, we provide data statistics and data analyses of the collected data to give a deeper understanding (§ 2.4). Finally, we discuss the data quality (§ 2.5).
@@ 2.1 Literature Book Mining
First, we extract all sentences from these books, and filter out too short or too long sentences, i.e., less than 10 words or more than 100 words, resulting in 577.6K literature sentences. Second, for each sentence, we use Qwen2.5-72B-Instruct Yang et al. (2024a) to judge whether the sentence involves similes or metaphors, and discard the sentences that do not contain any ones. Third, for the remaining sentences, we let Qwen2.5-72B-Instruct literally translate them to Chinese, and then judge whether the translation satisfies native Chinese people. If the answer is negative, the corresponding sentence will be reserved, and regarded as “suitable to translate via long thought”. For prompt details, please refer to Appendix A.1. Consequently, we collect 63K (out of 577.6K) literature sentences involving similes or metaphors whose literal translations have flaws, called pre-collected sentences.
@@ 2.2 Multi-Agent Framework
For each pre-collected sentence (denoted as $s$), we design a multi-agent framework to translate it via long thought. As shown in Figure 1, our multi-agent framework includes three agents: a translator, an advisor, and an evaluator, each of which use Qwen2.5-72B-Instruct as the backbone. The synthetic process is illustrated as follows:
@@
(3) Translation Refine Loop. In the refine loop, three agents work together to refine the translation iteratively. In each iteration step $k$ (start from $k=1$), the advisor first evaluates the translation in the previous step, i.e., $t^{k-1}$, and provides detailed feedback $f^{k-1}$ for polishing it. Then, the evaluator gives an overall score of $t^{k-1}$ conditioned on both pre-defined scoring criteria and $f^{k-1}$, and the score is denoted as $s^{k-1}$. In the last of the iteration step, the translator takes its previous translation $t^{k-1}$, the corresponding feedback $f^{k-1}$ and overall score $s^{k-1}$ into account to provide a new translation $t^{k}$. The translation refine loop will stop when the overall score reaches a pre-defined threshold or the number of iteration steps meets the maximum. For prompt details of the translator, advisor and evaluator, please refer to Appendix A.2.
@@ 2.3 Long Thought Reformulation
where $1\leq r_{1}<r_{2}<...<r_{n}\leq m$, and $n$ is the number of remaining steps. If $n<3$, we will discard the whole sample, i.e., $\mathcal{P}(s)$.
For the remaining samples, we follow Qin et al. (2024), and leverage GPT-4o OpenAI (2024a) to modify and polish $\mathcal{P}^{{}^{\prime}}(s)$ into a self-reflection description (the used prompt is provided in Appendix A.3). Finally, we obtain 22,264 MT samples with long thought. Figure 2 gives an example sample to illustrate the synthetic results.
@@
DRT data ${}_{\text{(training)}}$ 19,264 37.25 527.64 44.67
DRT data ${}_{\text{(validation)}}$ 1,000 37.43 531.36 44.98
DRT data ${}_{\text{(testing)}}$ 2,000 37.19 525.44 44.70
Table 1: The number of samples and average token-level length of query, thought and output. “Query” and “Output” in DRT data mean the source sentences and the translated outputs, respectively.
@@
Figure 3: The distribution of the number of refinement steps in DRT data.
2.4 Data Statistics and Data Analyses
Figure 4: Trends in average scores (provided by the evaluator agent) over the refinement steps. The trends for samples with three, four, five, and six refinement steps are illustrated in (a), (b), (c), and (d), respectively.
We split the collected 22,264 samples into training, validation and testing sets with 19,264, 1,000 and 2,000 samples, respectively. Table 1 shows the data statistics of DRT data and previous O1-like data. For Marco-O1 CoT data Zhao et al. (2024), since it is not fully released, we use its demo data to calculate the data statistics. [^4] https://github.com/AIDC-AI/Marco-o1 As we can see, the average number of tokens in our synthesized thought reaches 500+ tokens, showing the long thought process in our data.
Refine Loop Analyses. Figure 3 shows the number of refinement steps in the DRT data, which ranges from 3 to 8 steps. We can find that most samples (73.22%) involve 3 refinement steps, while only one sample involves 8 steps. Furthermore, to provide a deeper understanding of the refinement process, we calculate the average edit distance before and after each refinement step. Specifically, the first three refinement steps cause 21.44, 13.16 and 10.90 character-level edit distance. This observation is consistent with intuition. As the refinement progresses, the magnitude of the modification gradually decreases. To further understand the improvement brought by the translation refine loop, we calculate the average overall scores (provided by the evaluator agent) along with each refinement step. As shown in Figure 4, as the number of refinement steps increases, the average score generally increases, demonstrating that the refine loop could iteratively increase the quality of translations.
Metric ACC. (%)
CometKiwi 56.0
Evaluator Agent (Qwen2.5-72B-Instruct) 92.5
Evaluator Agent (GPT-4o) 93.5
Table 2: Accuracy of automatic metrics for translation quality estimation (ACC.: accuracy).
2.5 Quality Analyses
The Effectiveness of the Evaluator Agent. Previous work has shown that the state-of-the-art LLMs can be used as evaluators for various text generation tasks Kocmi and Federmann (2023); Wang et al. (2023); Li et al. (2024). To figure out the effectiveness of our evaluator agent, we randomly select 200 source sentences from DRT data, and for each of them, we further select its two translations as well as scores (provided by the evaluator agent) during refinement. We next employ human annotators to compare the two translations of each source sentence, and judge which translation is better, or two translations are similar in quality (annotation details can be found in Appendix B). After obtaining the quality labels, we calculate the accuracy of the evaluator agent according to its evaluation score. For comparison, we also calculate the accuracy of CometKiwi Rei et al. (2022) and GPT-4o evaluator agent. As shown in Table 2, our evaluator agent achieves a high accuracy (92.5%), demonstrating its effectiveness in evaluating literature translation quality. Besides, the widely-used CometKiwi metric only achieves 56.0% accuracy. Thought CometKiwi is powerful in the general domain (e.g., news) Kocmi and Federmann (2023), its effectiveness in the literature domain is limited and unreliable, which is also pointed out by Karpinska and Iyyer (2023). Furthermore, the GPT-4o evaluator agent slightly outperforms the origin evaluator agent (with Qwen2.5-72B-Instruct backbone). Considering the tradeoff between cost and effectiveness, we finally decide to use Qwen2.5-72B-Instruct as our evaluator agent.
Translation Quality. Based on the effectiveness of the evaluator agent and the observation that evaluation scores of final translations typically reach 90.0 (c.f., Figure 4), we can ensure a high level of translation quality in the constructed data. According to the pre-defined scoring criteria of the evaluator agent (c.f., Appendix A.2), a score of 90.0 indicates excellent translations.
@@ 2.5 Quality Analyses
GEA GRF CometKiwi GRB BLEU CometScore
@@
Llama-3.1-8B-Instruct 59.58 79.25 70.14 73.30 18.55 74.58
Qwen2.5-7B-Instruct 66.21 81.53 70.36 77.92 27.02 76.78
Qwen2.5-14B-Instruct 70.86 84.74 72.01 80.85 30.23 78.84
Marco-o1-7B 64.24 82.41 71.62 77.50 29.48 77.41
QwQ-32B-preview 75.50 86.31 71.48 83.08 27.46 78.68
DeepSeek-R1-Distill-Llama-8B 56.89 76.31 67.13 69.49 15.83 71.82
DeepSeek-R1-Distill-Qwen-7B 43.66 65.16 63.49 58.13 10.99 69.21
DeepSeek-R1-Distill-Qwen-14B 70.64 83.92 71.01 80.29 25.55 77.66
DeepSeek-R1-Distill-Qwen-32B 71.88 84.78 71.93 81.59 29.36 78.93
SFT LLMs (w/o CoT)
Llama-3.1-8B-SFT 69.33 84.10 70.25 80.18 30.03 78.26
Qwen2.5-7B-SFT 72.29 85.06 71.03 81.72 35.44 80.10
Qwen2.5-14B-SFT 74.53 85.66 72.08 83.08 37.63 80.82
DRT
DRT-8B (Backbone: ${}_{\text{(Backbone: Llama-3.1-8B-Instruct)}}$ 69.65 84.49 70.85 80.80 32.67 78.81
DRT-7B (Backbone: ${}_{\text{(Backbone: Qwen2.5-7B-Instruct)}}$ 75.05 85.57 71.78 82.38 35.54 80.19
DRT-14B (Backbone: ${}_{\text{(Backbone: Qwen2.5-14B-Instruct)}}$ 77.41 87.19 72.11 83.20 36.46 80.64
Table 3: Experimental results on literature translation. The bold and the underline denote the best and second-best performances, respectively. $\dagger$” and“$\ddagger$” denote statistically significant better than the corresponding SFT LLMs (w/o CoT) with t-test p < 0.01 and 0.05, respectively.
@@ 3.1 Experimental Setups
Metrics. Following previous work, we adopt “BLEU” Papineni et al. (2002), “CometKiwi” and “CometScore” Rei et al. (2022) to evaluate the model translations. Among them, BLEU evaluates n-grams overlap between model translations and references, while CometScore evaluates the semantic similarity of model translations against references. CometKiwi uses a language model to judge whether a model translation conveys the semantics of the source sentence.
@@
As pointed out by Karpinska and Iyyer (2023), BLEU and COMET may be ineffective for evaluating literature translation. Meanwhile, recent studies also show the strong ability of LLMs in NLP evaluation Li et al. (2024). Therefore, we use evaluators implemented using GPT-4o in reference-based and reference-free styles, which we refer to as “GRB” and “GRF”, respectively. The evaluation prompts borrow from Kocmi and Federmann (2023), and are illustrated in Appendix C. Furthermore, as demonstrated in § 2.4, the GPT-4o evaluator agent achieves great accuracy in literature translation. We also leverage it as the evaluation metric in experiments, which is referred to as “GEA”. Since GRB, GRF and GEA need the API costs, we randomly select 400 samples to conduct evaluation.
@@
Backbones. We adopt the following three LLMs as the backbones of our DRT: Llama-3.1-8B-Instruct Dubey et al. (2024), Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct Yang et al. (2024b). All model checkpoints are publicly available.
For evaluation toolkits and the implementation details of all models, please refer to Appendix D.
3.2 Comparison Models
Vanilla LLMs. We leverage vanilla Llama-3.1-8B-Instruct, Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct Yang et al. (2024a) as the comparison models. Besides, six O1-like LLMs are also conducted as baselines: Marco-o1-7B Zhao et al. (2024), QwQ-32B-preview Qwen (2024), DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Qwen-14B and DeepSeek-R1-Distill-Qwen-32B Guo et al. (2025).
SFT LLMs (w/o CoT). We also fine-tune LLMs with only paired sentences of DRT training data (without thought). This setting allows LLMs to learn the mapping from source literature sentences to the corresponding Chinese translations directly. We denote the fine-tuned LLMs as Llama-3.1-8B-SFT, Qwen2.5-7B-SFT and Qwen2.5-14B-SFT, serving as strong baselines in the experiments.
3.3 Main Results
Table 3 shows the experimental results, we analyze the performance from the following aspects:
SFT LLMs (w/o CoT) vs. Vanilla LLMs. After instruction tuning on the paired sentences of our training data, SFT LLMs (w/o CoT) significantly outperform the corresponding vanilla LLMs. For example, Llama-3.1-8B-SFT outperforms Llama-3.1-8B-Instruct by 9.75 GEA, 4.85 GRF and 6.88 GRB. Qwen2.5-7B-SFT outperforms Qwen2.5-7B-Instruct by 6.08 GEA, 3.53 GRF and 3.80 GRB. This finding demonstrates the effectiveness of our multi-agent framework and the quality of the synthesized translation. Please also note that the final translations are synthesized by Qwen2.5-72B-Instruct, indicating that we can leverage off-the-shelf open-source LLMs to collect high-quality literation translation data. And the data could help smaller LLMs (such as 7B and 14B ones) to boost their literature translation skills.
DRT vs. Vanilla LLMs. After fine-tuning on the long-thought MT training data, our DRT-series LLMs also significantly outperform the corresponding vanilla backbones. Particularly, DRT-14B outperforms QwQ-32B-preview and DeepSeek-R1-Distill-Qwen-32B in terms of all metrics, showing its effectiveness in literature MT.
DRT vs. SFT LLMs (w/o CoT). Using Llama-3.1-8B-Instruct and Qwen2.5-7B-Instruct as backbones, LLMs tuned with long thought achieve better performance than those tuned without long thought in terms of all metrics. For example, DRT-7B outperforms Qwen2.5-7B-SFT by 2.76 GEA, 0.51 GRF, 0.75 CometKiwi, 0.66 GRB, 0.10 BLEU and 0.09 CometScore. When using Qwen2.5-14B-Instruct as the backbone, we find that DRT-14B outperforms Qwen2.5-14B-SFT in terms of GEA, GRF, CometKiwi and GRB, but underperforms in terms of BLEU and CometScore. In detail, BLEU and CometScore evaluate the translations from the perspective of similarity between model translations and golden references. We conjecture that the higher BLEU and CometScore performance of Qwen2.5-14B-SFT is due to the model’s ability to quickly learn domain-specific translations through tuning without long thoughts, allowing it to adapt to the literature translation more straightforwardly. However, training without long thoughts might lead the model to a sub-optimal solution, like learning shortcuts. When adopting evaluation metrics that are not significantly dependent on the golden references (i.e., GEA, GRF, CometKiwi and GRB), DRT-14B shows its superior performance. Note that although GRB is a reference-based metric, it does not assess the model translations simply based on how similar they are to the golden references.
DRT vs. Commercial LLMs. To give a deeper understanding of our DRT models’ performance, we further compare DRT models with GPT-4o OpenAI (2024a) and o1-preview OpenAI (2024b). The experimental results and corresponding analyses are provided in Appendix E.
Model Flu. Sem. Lit.
Qwen2.5-14B-Instruct -0.353 -0.363 -0.442
QwQ-32B-Preview -0.063 0.022 -0.007
Qwen2.5-14B-SFT 0.103 0.108 0.087
DRT-14B 0.313 0.233 0.362
Table 4: Human evaluation results in terms of fluency, semantic accuracy and literary quality.
3.4 Human Evaluation
We conduct human evaluation to further evaluate the performance of DRT-14B and strong baselines (Qwen2.5-14B-Instruct, QwQ-32B-Preview and Qwen2.5-14B-SFT). We randomly select 200 samples from our test set, and employ three human evaluators with high levels of fluency in English and Chinese to assess the generated translations from three aspects: fluency (Flu.), semantic accuracy (Sem.) and literary quality (Lit.). Following the Best-Worst Scaling method Kiritchenko and Mohammad (2017), evaluators are asked to select the best and the worst generated translation on each aspect. The result scores are calculated based on the percentage of times each model is selected as best minus the times it is selected as worst. Thus, the final scores should range from -1 (worst) to 1 (best). As shown in Table 4, DRT-14B significantly outperforms these strong baselines, especially in the aspect of literary quality. These results demonstrate the superiority of our DRT models. The Fleiss’ Kappa scores Fleiss (1971) of Flu., Sem. and Lit. are 0.75, 0.69 and 0.85, respectively, indicating a good inter-agreement among evaluators.
3.5 Inference Time Analysis
During evaluating LLMs’ literature translation performance on our test set, we leverage vLLM to accelerate the model generation. A single NVIDIA A100 GPU (40G) is used to deploy each LLM. As shown in Figure 5, the average time costs of DRT models are significantly higher than LLMs (w/o CoT). This is because DRT models should first generate the long thought and then provide the final translation, thus needing more inference time ($\times$11.9~13.9). This also indicates that the O1-like LLMs may not be applicable to some scenarios with high real-time requirements.
Figure 5: Time cost during inference on the testing set.
Source Sentence Qwen2.5-14B-Instruct QwQ-32B-Preview Qwen2.5-14B-SFT DRT-14B